# Faster Convergence of Riemannian Stochastic Gradient Descent with Increasing Batch Size

### v5

@@ 2 Preliminaries
Iterative methods on Riemannian manifold are defined by $x_{t+1}=R_{x_{t}}(\eta_{t}d_{t})$ generally. Exponential maps are often used as retractions. The following assumption plays a central role in Lemma 3.1. {assumption} [Retraction Smoothness] Let $f\colon\mathcal{M}\rightarrow\mathbb{R}$ be a smooth map. Then there exists $L_{r}>0$, such that $^{\forall}x\in\mathcal{M},^{\forall}v\in T_{x}\mathcal{M}$, $f(R_{x}(v))\leq f(x)+\langle\mathrm{grad}f(x),v\rangle_{x}+\frac{L_{r}}{2}\|v\|_{x}^{2}$. In the Euclidean space setting, $L$-smoothness implies a property similar to retraction smoothness. The property corresponding to $L$-smoothness in Euclidean space is defined for $f\colon\mathcal{M}\rightarrow\mathbb{R}$ as $^{\exists}L>0,^{\forall}x,y\in\mathcal{M}\colon\|\mathrm{grad}f(x)-\Gamma^{x}_{y}\mathrm{grad}f(y)\|_{x}\leq L_{r}d(x,y)$, where $\Gamma$ is the parallel transport from $T_{y}\mathcal{M}$ to $T_{x}\mathcal{M}$, and $d(\cdot,\cdot)$ is the Riemannian distance. This condition is sufficient for Assumption 2.1 with $R\coloneq\mathrm{Exp}$ (see Boumal (2023, Corollary 10.54)). This case is frequently used (e.g., Zhang and Sra (2016); Criscitiello and Boumal (2023); Kim and Yang (2022); Liu et al. (2017)). Other sufficient conditions for Assumption 2.1 were identified by Kasai et al. (2018, Lemma 3.5) and Sakai and Iiduka (2025, Proposition 3.2). Now, we consider the empirical risk minimization problem such that
@@
where each $f_{j}\colon\mathcal{M}\rightarrow\mathbb{R}$ is smooth and lower bounded. Therefore, $f$ is also smooth and lower bounded. This assumption is often made for both Euclidean space and Riemannian space. The lower boundedness of $f$ is essential for analyses using optimization theory because unbounded $f$ may not have optimizers. We denote an optimal value of $f$ by $f^{\star}$ and let $N$ denote the size of the dataset. In many machine learning tasks, either the dimension of model parameters $x$ or the size of dataset $N$ is large. Hence, we use the following minibatch gradient to efficiently approximate the gradient of $f$: $\mathrm{grad}f_{B}(x)\coloneq\frac{1}{b}\sum_{j=1}^{b}\mathrm{grad}f_{\xi_{j}}(x)$, where $B$ represents a minibatch with size $b$, and $(\xi_{j})_{j=1}^{b}$ is a sequence of $\{1,\cdots,N\}$-valued i.i.d. random variables distributed as $\Pi$. That is, the gradient of $f_{B}(x)=\frac{1}{b}\sum_{j=1}^{b}f_{\xi_{j}}(x)$ is used instead of $\mathrm{grad}f(x)$. The following assumption is reasonable given this situation. {assumption} [Bounded Variance Estimator] The stochastic gradient given by a distribution $\Pi$ is an unbiased estimator of the full gradient and has bounded variance: (I) $\forall x\in\mathcal{M}\colon\mathbb{E}_{\xi\sim\Pi}\left[\mathrm{grad}f_{\xi}(x)\right]=\mathrm{grad}f(x)$; (II) $\exists\sigma>0,\forall x\in\mathcal{M}\colon\mathbb{V}_{\xi\sim\Pi}(\mathrm{grad}f_{\xi}(x))\leq\sigma^{2}$. For example, if $f$ is lower bounded, satisfies Assumption 2.1, and $\Pi$ is taken to be the uniform distribution over $\{1,\ldots,N\}$, then Assumption 2 holds. RSGD is defined by $x_{t+1}=R_{x_{t}}(-\eta_{t}\mathrm{grad}f_{B_{t}}(x_{t})).$ We assume that the sequence $(x_{t})_{t}$ generated by RSGD is independent of the random variables $(\xi_{j,t})_{j,t}$ used in the minibatch gradient. Note that the BS $(b_{t})_{t}$ may vary at each iteration.
@@ 3 Convergence Analysis
Let $(x_{t})_{t}$ be a sequence generated by RSGD and let $\eta_{\max}>0$. Consider a positive-valued sequence $(\eta_{t})_{t}$ such that $\eta_{t}\in[0,\eta_{\max}]\subset[0,\frac{2}{L_{r}})$. Then, under Assumptions 2.1 and 2, we obtain
@@ C.1 Gradient Norm versus SFO Complexity for Our Proposed] Optimization Problem
Appendix D Proof of our Lemma and Theorems
We define $\mathcal{F}_{t}\coloneq\sigma[x_{0},(\xi_{j,0})_{j},(\xi_{j,1})_{j},\ldots,(\xi_{j,t-1})_{j}]$, where $x_{0}$ is the initial point of RSGD. For a measurable function $f$, $\mathbb{E}[f|\mathcal{F}_{t}]$ denotes the conditional expectation with respect to $\mathcal{F}_{t}$. Note that, from Assumption 2, $\mathbb{E}_{\xi\sim\Pi}\left[\|\mathrm{grad}f_{\xi}(x)-\mathrm{grad}f(x)\|_{x}^{2}\right]=\mathbb{V}_{\xi\sim\Pi}(\mathrm{grad}f_{\xi}(x))\leq\sigma^{2}$ holds. The following result serves as a preliminary for Lemma 3.1.
@@ Appendix D Proof of our Lemma and Theorems
Let $(x_{t})_{t}$ be a sequence generated by RSGD and $(\eta_{t})_{t}$ be a positive-valued sequence. Then, under Assumptions 2.1 and 2, we obtain

### v6

@@ 2 Preliminaries
Iterative methods on Riemannian manifold are defined by $x_{t+1}=R_{x_{t}}(\eta_{t}d_{t})$ generally. Exponential maps are often used as retractions. The following assumption plays a central role in Lemma 3.1.
Assumption 2.2 (Retraction Smoothness)
Let $f\colon\mathcal{M}\rightarrow\mathbb{R}$ be a smooth map. Then there exists $L_{r}>0$, such that $^{\forall}x\in\mathcal{M},^{\forall}v\in T_{x}\mathcal{M}$, $f(R_{x}(v))\leq f(x)+\langle\mathrm{grad}f(x),v\rangle_{x}+\frac{L_{r}}{2}\|v\|_{x}^{2}$.
In the Euclidean space setting, $L$-smoothness implies a property similar to retraction smoothness. The property corresponding to $L$-smoothness in Euclidean space is defined for $f\colon\mathcal{M}\rightarrow\mathbb{R}$ as $^{\exists}L>0,^{\forall}x,y\in\mathcal{M}\colon\|\mathrm{grad}f(x)-\Gamma^{x}_{y}\mathrm{grad}f(y)\|_{x}\leq L_{r}d(x,y)$, where $\Gamma$ is the parallel transport from $T_{y}\mathcal{M}$ to $T_{x}\mathcal{M}$, and $d(\cdot,\cdot)$ is the Riemannian distance. This condition is sufficient for Assumption 2.2 with $R\coloneq\mathrm{Exp}$ (see Boumal (2023, Corollary 10.54)). This case is frequently used (e.g., Zhang and Sra (2016); Criscitiello and Boumal (2023); Kim and Yang (2022); Liu et al. (2017)). Other sufficient conditions for Assumption 2.2 were identified by Kasai et al. (2018, Lemma 3.5) and Sakai and Iiduka (2025, Proposition 3.2). Now, we consider the empirical risk minimization problem such that
@@
where each $f_{j}\colon\mathcal{M}\rightarrow\mathbb{R}$ is smooth and lower bounded. Therefore, $f$ is also smooth and lower bounded. This assumption is often made for both Euclidean space and Riemannian space. The lower boundedness of $f$ is essential for analyses using optimization theory because unbounded $f$ may not have optimizers. We denote an optimal value of $f$ by $f^{\star}$ and let $N$ denote the size of the dataset. In many machine learning tasks, either the dimension of model parameters $x$ or the size of dataset $N$ is large. Hence, we use the following minibatch gradient to efficiently approximate the gradient of $f$: $\mathrm{grad}f_{B}(x)\coloneq\frac{1}{b}\sum_{j=1}^{b}\mathrm{grad}f_{\xi_{j}}(x)$, where $B$ represents a minibatch with size $b$, and $(\xi_{j})_{j=1}^{b}$ is a sequence of $\{1,\cdots,N\}$-valued i.i.d. random variables distributed as $\Pi$. That is, the gradient of $f_{B}(x)=\frac{1}{b}\sum_{j=1}^{b}f_{\xi_{j}}(x)$ is used instead of $\mathrm{grad}f(x)$. The following assumption is reasonable given this situation.
Assumption 2.3 (Bounded Variance Estimator)
The stochastic gradient given by a distribution $\Pi$ is an unbiased estimator of the full gradient and has bounded variance: (I) $\forall x\in\mathcal{M}\colon\mathbb{E}_{\xi\sim\Pi}\left[\mathrm{grad}f_{\xi}(x)\right]=\mathrm{grad}f(x)$; (II) $\exists\sigma>0,\forall x\in\mathcal{M}\colon\mathbb{V}_{\xi\sim\Pi}(\mathrm{grad}f_{\xi}(x))\leq\sigma^{2}$.
For example, if $f$ is lower bounded, satisfies Assumption 2.2, and $\Pi$ is taken to be the uniform distribution over $\{1,\ldots,N\}$, then Assumption 2.3 holds. RSGD is defined by $x_{t+1}=R_{x_{t}}(-\eta_{t}\mathrm{grad}f_{B_{t}}(x_{t}))$, and is a generalization of Euclidean SGD. Because $(\xi_{i,t})_{i,t}$ are i.i.d. samples, $\boldsymbol{\xi}_{T}\coloneq(\xi_{1,T},\cdots,\xi_{b_{T},T})^{\top}$ is independent of $(x_{t})_{t=0}^{T}$. Note that the BS $(b_{t})_{t}$ may vary at each iteration. The detailed definitions of $\mathbb{E}_{\xi\sim\Pi},\mathbb{V}_{\xi\sim\Pi},$ and $\mathbb{E}$ can be found in Appendix D.
@@ 3 Convergence Analysis
Let $(x_{t})_{t}$ be a sequence generated by RSGD and let $\eta_{\max}>0$. Consider a positive-valued sequence $(\eta_{t})_{t}$ such that $\eta_{t}\in[0,\eta_{\max}]\subset[0,\frac{2}{L_{r}})$. Then, under Assumptions 2.2 and 2.3, we obtain
@@ C.1 Gradient Norm versus SFO Complexity for Our Proposed] Optimization Problem
Appendix D Proofs of our Lemma and Theorems
Recall that $\Pi$ is a $\{1,\cdots,N\}$-valued probability distribution. Thus, we consider the probability space $(\{1,\cdots,N\},\mathfrak{P}(\{1,\cdots,N\}),\Pi)$. For a random variable $\xi$ distributed as $\Pi$ and a function $g:\{1,\cdots,N\}\times\mathcal{M}\ni(\xi,x)\mapsto g_{\xi}(x)\in\mathbb{R}$, we define $\mathbb{E}_{\xi}[g_{\xi}(x)]\coloneq\mathbb{E}_{\xi\sim\Pi}[g_{\xi}(x)]\coloneq\int_{\{1,\cdots,N\}}g_{\xi}(x)\Pi(d\xi)=\sum_{j=1}^{N}g_{j}(x)\Pi(\{j\})$. Hence, the variance is defined as $\mathbb{V}_{\xi\sim\Pi}(g_{\xi}(x))\coloneq\mathbb{E}_{\xi\sim\Pi}[\left\|g_{\xi}(x)-\mathbb{E}_{\xi\sim\Pi}[g_{\xi}(x)]\right\|^{2}]$. Let $(\xi_{i,t})_{i=1}^{N}$ be a sequence of i.i.d. random variables distributed as $\Pi$, let $\boldsymbol{\xi}_{t}\coloneq(\xi_{1,t},\cdots,\xi_{b_{t},t})^{\top}\in\{1,\cdots,N\}^{b_{t}}$, let $g:\{1,\cdots,N\}^{b_{t}}\times\mathcal{M}\ni(\boldsymbol{\xi},x)\mapsto g_{\boldsymbol{\xi}}(x)\in\mathbb{R}$ be a function, and let $x_{t}$ be a point at the $t$-th iteration generated by the updating rule of RSGD. We can introduce a natural extension of the expectation notation to a multivariate random variable:
$\mathbb{E}_{\boldsymbol{\xi}_{t}}[g_{\boldsymbol{\xi}_{t}}(x_{t})]$ $\coloneq\mathbb{E}_{\boldsymbol{\xi}_{t}\sim\Pi^{b_{t}}}[g_{\boldsymbol{\xi}_{t}}(x_{t})]\coloneq\int_{\{1,\cdots,N\}^{b_{t}}}g_{\boldsymbol{\xi}_{t}}(x_{t})\Pi^{b_{t}}(d\boldsymbol{\xi}_{t})$
$=\sum_{j_{1}=1}^{N}\cdots\sum_{j_{b_{t}}=1}^{N}g_{(j_{1},\cdots,j_{b_{t}})^{\top}}(x_{t})\Pi(\{j_{1}\})\cdots\Pi(\{j_{b_{t}}\})$
$=\mathbb{E}_{\xi_{1,t}}\mathbb{E}_{\xi_{2,t}}\cdots\mathbb{E}_{\xi_{b_{t},t}}[g_{\boldsymbol{\xi}_{t}}(x_{t})].$
For $\boldsymbol{\xi}_{1},\cdots,\boldsymbol{\xi}_{t}$, a total expectation is defined as $\mathbb{E}\coloneq\mathbb{E}_{\boldsymbol{\xi}_{1}}\cdots\mathbb{E}_{\boldsymbol{\xi}_{t}}$. Note that, from Assumption 2.3, $\mathbb{E}_{\xi}\left[\|\mathrm{grad}f_{\xi}(x)-\mathrm{grad}f(x)\|_{x}^{2}\right]=\mathbb{V}_{\xi}(\mathrm{grad}f_{\xi}(x))\leq\sigma^{2}$ holds, which implies $\mathbb{E}_{\boldsymbol{\xi}}\left[\|\mathrm{grad}f_{\boldsymbol{\xi}}(x)-\mathrm{grad}f(x)\|_{x}^{2}\right]\leq\frac{\sigma^{2}}{b_{t}}$. The following result serves as a preliminary for Lemma 3.1.
@@ Appendix D Proofs of our Lemma and Theorems
Let $(x_{t})_{t}$ be a sequence generated by RSGD and $(\eta_{t})_{t}$ be a positive-valued sequence. Then, under Assumptions 2.2 and 2.3, we obtain